Withoust consultation , each of four organizations announced a one-daY convention to be held during june. Find the probability that at least two organizations specify the same day for their convention?
Answer by stanbon(75887)

Withoust consultation , each of four organizations announced a one-daY convention to be held during june. Find the probability that at least two organizations specify the same day for their convention?
----
Binomial Problem with n=4 and p = 1/7
-----
P(2<= x <=4)
= 1 - P(0<= x <=1)
= 1 - binomcdf(4,1/7,1)
= 0.1004
================
Cheers,
Stan H.
================